OpenStudy (anonymous):

Calculate the sum of 1/4 and 1/7

OpenStudy (anonymous):

any ideas?

OpenStudy (anonymous):

None

OpenStudy (anonymous):

okay so convert those fractions to decimals

OpenStudy (anonymous):

then try and find the sum of that

OpenStudy (anonymous):

.39

OpenStudy (anonymous):

now make that a fraction

OpenStudy (anonymous):

1/4 + 1/7= [1 *7]/[4 *7] + [1 *4]/7[1 *4] = 7/28 + 4/28 = [7+4] /28 = 11/28